the sum of the present ages of tahir and masud is 18 years. Three years ago tahir was three times as old as masud. find their present ages.
Answer by scott8148(6628)

t + m = 18

t - 3 = 3(m - 3)

subtracting equations (to eliminate t) ___ -3 - m = 3(m - 3) - 18 ___ -3 - m = 3m - 27 ___ 24 = 4m

solve for m, then substitute back to find t